solve
x+2y=0
x-2y=12

Answer by checkley71(8403)

X+2Y=0
X-2Y=12 BY ADDING THESE TWO EQUATIONS WE GET
----------
2X=12 OR X=12/2 OR X=6 THEN 6+2Y=0 OR 2Y=-6 OR Y=-6/2 OR Y=-3